# Existence of Decreasing Nambu Solutions to the Rainbow–Ladder Gap Equation of QCD by Cone Compression

### v1

@@ 1  Introduction
Establishing the existence of solutions to the gap equation in QCD is not easy; however, the spectrum of solutions has been shown to be rich[12] and at finite temperature, the chiral symmetry restoration transition was suggested to be of second-order[4] due to the structure of the fermion propagator which was later verified in numerical simulations[1]. [^2] However, at very small or large quark masses, this is expected to change[3]. A related phase transition is the emergence of Nambu solutions as the coupling is increased past the dynamical chiral symmetry breaking (DCSB) critical point which we study here at $T=0$.
@@
For a primer on the gap equation for mathematicians and fixed-point theory for physicists, see Appendix A,B. We analyze the solution set of the gap equation after the renormalization scale has been taken to infinity in the equation, see the Appendix for details. Solutions clearly also satisfy a finite renormalization scale equation for some current quark mass parameter but the reverse is not guaranteed. In the case of the gap equation, the current quark mass parameter disappears in the limit of infinite renormalization scale and so any solutions which rely on it to reach an eigenvalue of 1 also disappear with it, namely the Wigner solutions. Solutions that set their own scale as the current quark mass parameter disappears, namely the Nambu solutions, survive and are guaranteed to exist for all $m\equiv\lim_{p\rightarrow\infty}(\log\frac{p}{\mu})^{\gamma_{m}}M(p)$ as we will prove for a popular model using the rainbow-ladder(RL) vertex[9].
@@
Existence results of non-linear integral equations $u=T(u)\equiv\int_{0}^{\infty}dkK(k,p)f(u(k),k)$, often aim to establish a fixed point between two opposite behaviors, say as $\|u\|\rightarrow 0$ and $\|u\|\rightarrow\infty$ with the supremum norm via the Krasnosel’skii-Guo Cone Compression Theorem[7, 8]. The trace of the gap equation takes exactly this form after sending the renormalization scale to $\infty$[10]. Another approach called Schauder Fixed Point Theorem can be used if $T(P)\subseteq P$ for some compact input space $P$ as we will see is the case for the trace of $\not${p} times the gap equation. Coupled systems with Krasnosel’skii type conditions in both input function have been analyzed in [6] and a mixed Krasnosel’skii-Schauder theorem has been developed in [5].
@@
Taking the trace of the Euclidean space rainbow ladder gap equation, we define $u(p)\equiv r(p)\frac{M(p)}{Z(p)}\equiv r(p)B(p)$ for some continuous and positive function $r(p)$ with $\lim_{p\rightarrow\infty}r(p)\sim\log^{\delta}\frac{p}{\mu}$ [^3] We write $\lim_{p\rightarrow p_{0}}A(p)\sim B(p)$ for $\lim_{p\rightarrow p_{0}}\frac{A(p)}{B(p)}=1$. for some $0<\delta<\gamma_{m}$ because it allows bounding the gap equation by an integrable kernel and a supremum or infimum over a bounded function. In the perturbative region, we show in the appendix that the solution is a sum of the expected $B_{+}(x)=\frac{c_{1}}{x^{\gamma_{m}}}$ with $x\equiv\log\frac{p}{\mu}$ and $B_{-}(x)=c_{2}x^{\gamma_{m}-1}e^{-2x}$. There, we prove that $c_{2}$ is always equal to a certain integral and we will show that $m\equiv c_{1}\geq 0$ is in fact free to choose for a certain guaranteed set of solutions. Consider the input space $P=\{B(x)\in C[(-\infty,\infty)]\ |\ B(x)\geq 0,\exists\gamma_{m}>0:\lim_{x\rightarrow\infty}\frac{\partial\log B}{\partial\log x}=-\gamma_{m}\}$. Since $r(p)$ is given, we can switch between $u$ and $B$ whenever convenient and $\lim_{x\rightarrow\infty}\frac{\partial\log B}{\partial\log x}=-\gamma_{m}$ implies $\lim_{x\rightarrow\infty}\frac{\partial\log T(B)}{\partial\log x}=-\gamma_{m}$ by D.1. We have
@@
with $q_{\mu}=p_{\mu}-k_{\mu}$ with $S^{3}$ the surface of the sphere that leaves $k^{2},p^{2}$ invariant. We first consider $Z(p)$ to be a known function that one can take to satisfy the gap equation if a solution exists. It should be almost everywhere continuous and converge to a perturbative tail: $Z(p)\rightarrow 1$ as $p\rightarrow\infty$. One can write $T=T^{Z}$ as the results will apply to one unchanging $Z(p)$. We extend our analysis to the coupled system of equations before the conclusion as Schauder’s theorem is required to deal with the equation for $Z(p)$. Notice $f(u,k)$ is positive and continuous and we require that $K(k,p)>0$ is continuous almost everywhere, $\int_{0}^{\infty}K(k,p)dk$ is continuous and $\mathcal{G}(q^{2})\sim\mathcal{G}_{pert}(q^{2})=\frac{4\pi^{2}\gamma_{m}}{q^{2}\log(\frac{q^{2}}{\mu^{2}})}$ as $q^{2}\rightarrow\infty$. [^4] QED also fulfills similar requirements if the Planck scale is taken as upper limit. We define a simplest extension $\mathcal{G}_{simp}(q^{2})=\frac{4\pi^{2}\gamma_{m}}{\max(q^{2},\mu^{2})\log(e+\frac{q^{2}}{\mu^{2}})}$ as a toy example.
@@ 2.3 Combining the bounds
$t(p)=\frac{1}{K_{r}}\int_{0}^{\infty}dk\frac{kZ^{2}(k)}{4\pi^{4}}\int_{S^{3}}\mathcal{G}(q^{2}(k,p))t(k)\equiv\frac{1}{K_{r}}T_{c}(t)$ (8)
@@ 3.1 Krasnosel’skii-Guo Theorem for $B(p)$
In order to utilize the Krasnosel’skii-Guo Theorem, we need to augment our space with a norm. Writing $\|u\|_{\infty}\equiv\sup_{p}u(p)$, consider therefore the Banach space $E$ of continuous functions on $p\in[0,\infty)$ such that $\lim_{p\rightarrow\infty}u(p)=0$ with norm $\|u\|=\|u\|_{\infty}+\sup_{\lambda>1,p>\mu}\log^{\gamma_{m}-\delta}(\lambda)u(\lambda p)$. The purpose of the second piece will soon become apparent but it can be checked that this is indeed a Banach space. Then $P_{\gamma_{m}}\subset P\subset E$ with $P_{\gamma_{m}}$ the positive functions with asymptotic corresponding to $\gamma_{m}$. Let $P_{\Delta}$ be the space corresponding to $\gamma_{m}=0$ and take one choice of $B=B_{0}(p)$ with asymptotic mass $m$. Clearly, $P_{\Delta}$ is a total cone in $E$ and for any $\in P_{\gamma_{m}}, BB$$B_{0}\in P_{\Delta}$ with $\|B_{0}\|_{\infty}<\epsilon$ for some small $\epsilon$. We denote the corresponding $u_{0}(p)=B_{0}(p)r(p)$ and $T_{\Delta}(u)\equiv T(u+u_{0})-u_{0}$.
$K(\cdot,p)$ is $L^{1}$-continuous by Scheffé’s lemma which gives equicontinuity and then compactness of $T_{\Delta}$ under the $\|u\|_{\infty}$ part of the norm via the Arzelà–Ascoli theorem. For the second part of the norm, the same follows for each $\lambda$ after a variable transformation $s=p\lambda$. To check that the norm cannot blow up as $\lambda\rightarrow\infty$, we use the asymptotic bound $\lim_{k\rightarrow\infty}u(k)=o(\frac{1}{\log^{\gamma_{m}-\delta}\frac{k}{\mu}})$ so that $\lim_{\lambda\rightarrow\infty}\log^{\gamma_{m}-\delta}(\lambda)u(\lambda p)=0$.
@@
$=\sup_{0<n<1}\lim_{\|u\|\rightarrow\infty}\sup_{0<n<1}\frac{n\frac{\gamma_{m}}{\delta}(1+\log^{\gamma_{m}-\delta}(\frac{\|u\|_{\infty}}{\mu})}{(1+\left(\frac{1}{1-n}\right)^{\frac{\gamma_{m}-\delta}{\delta}}\log^{\gamma_{m}-\delta}(\frac{\|u\|_{\infty}}{\mu}))}$ (10)
@@ 3.2 Schauder Theorem for $Z(p)$ and coupled equations
The gap equation for $Z(p)$ is given by Eqn.32. We require $K_{Z}(k,p)$ to be $L^{1}$-continuous. Then uniform boundedness follows by the reciprocal nature of the gap equation for $A(p)$ for an input space $\{A(p)\in C[[0,\infty)]\ |\ A(p)>A_{-}(p)>0\}$ and therefore compactness by the Arzelà–Ascoli theorem. If, as in the models in [9], the kernel $K_{Z}(k,p)$ is negative for $k<k^{*}(p)$ and positive for $k>k^{*}(p)$ for each $p$ and we define $T_{\Delta}:P^{r,R}_{\Delta,*}\rightarrow P_{\Delta}$ as before with $u(p)\equiv M(p)r(p)$ and $r(p)$ increasing, then because $M(p)$ is decreasing, the finite mass term will suppress the negative part of the kernel $K^{-}_{Z}$ more than the positive part and push $T(A)$ from $T(A_{0})$ in the direction of $A=1$ so that for $A(p)\geq\phi_{0}(p)\equiv\min[1,A_{0}(p)]$, we have $\phi_{0}(p)\leq T_{Z}(A)(p)\leq T^{+}_{Z}(\phi_{0})(p)$ for finite $M(p)$ where $T^{+}_{Z}$ is the positive part of the kernel (likewise for $T^{-}_{Z}$ so that $T_{Z}=T_{Z}^{+}+T_{Z}^{-}$). Therefore, we can use a combined Krasnosel’skii-Schauder theorem that has been derived for coupled systems[5] as long as $\sup_{r}K_{r}>1$ for $A=T_{Z}^{+}(\phi_{0})$. In fact, for this value a solution for $A(p)$ to $M=0$ can also be guaranteed by noticing that if $A_{-}(p)$ can be found such that $A(p)>A_{-}(p)$ and $A_{--}(p)\equiv((T_{Z}^{+})^{2}+T^{-}_{Z})(A_{-})>A_{-}(p)$, then $A_{--}(p)<T(A)(p)<T^{+}_{Z}(A_{-})(p)$, [^9] Sufficient condition for $A_{-}(p)$ to be a lower bound and therefore $A_{+}=T^{+}_{Z}(A_{-})$ to be an upper bound is $T^{-}_{Z}(A_{-})+T_{Z}^{+}(A_{+})>A_{-}$ for all $p$. guaranteeing a solution for $M=0$ by Schauder’s theorem and to the coupled system as long as $\sup_{r}K_{r}>1$ for $A(p)=T^{+}_{Z}(\min[1,A_{-}(p)])$. Therefore, we have $(T_{\Delta},T_{Z}):P^{r,R}_{\Delta,*}\times P_{Z}\rightarrow P_{\Delta}\times P_{Z}$ with $T_{\Delta}$ allowing us to use the Krasnosel’skii-Guo Theorem as before [^10] The version used in [5] uses a more general formulation due to Benjamin which however includes all cases where the Krasnosel’skii-Guo Theorem holds. and $P_{Z}=\{A\in C[[0,\infty)]\ |\ A_{-}(p)\leq A(p)\leq T_{Z}^{+}(A_{-})(p)\}$ a closed, convex and bounded set under sup-norm which is closed under $T_{Z}$, therefore amenable to Schauder’s Theorem. Therefore, a positive, continuous solution pair $(u(p),A(p))$ with $u(p)$ decreasing must exist for every value of $m$ by the combined theorem. We require $\sup_{r^{*}}K_{r}>1$ for $A(p)=T^{+}_{Z}(\min[1,A_{-}(p)])$ and $r^{*}(p)$ increasing so that $r(p)$ can be chosen to be increasing and $M(p)$ is decreasing.
@@ Conclusion
Considering a more general vertex, for example the Ball-Chui vertex[10, 2], we encounter serious difficulties with this approach due to the fact that a finite-difference derivative term $A^{\prime}(p)$ appears in the equation for $B(p)$ which is not constrained to have any sign or magnitude by Schauder’s Theorem.
Part of the issue with quantum field theory is that the infinite tower of Dyson-Schwinger equations may not have a guaranteed solution space. A corollary to the results in this paper is that for the gap equation in QCD, the contractive condition at $\|u\|\rightarrow\infty$ for the Krasnosel’skii-Guo theorem is satisfied for any $m$.
@@ Appendix A Gap equation primer for mathematicians
The gap equation is a 2-component coupled integral equation for the quark mass function $M(p)$ and field renormalization function $Z(p)$ as a function of 4-momentum $p_{\mu}$ with $p\equiv\sqrt{-p^{2}}$ in terms of the gluon kernel $\mathcal{G}(q^{2})$ and vertex $\Gamma_{\mu}(p,k,q)$ with $k_{\mu}$ and $q_{\mu}=p_{\mu}-k_{\mu}$ loop momenta. The simplification of the rainbow-ladder vertex amounts to $\Gamma_{\mu}(p,k,q)=\gamma_{\mu}$. After rotating to the Euclidean metric, letting $A(p)\equiv\frac{1}{Z(p)}$ and $B(p)\equiv\frac{M(p)}{Z(p)}$, we have the operators

### v2

@@ 1  Introduction
Establishing the existence of solutions to the gap equation in QCD is not easy; however, the spectrum of solutions has been shown to be rich[13] and at finite temperature, the chiral symmetry restoration transition was suggested to be of second-order[5] due to the structure of the fermion propagator which was later verified in numerical simulations[2]. [^2] However, at very small or large quark masses, this is expected to change[4]. A related phase transition is the emergence of Nambu solutions as the coupling is increased past the dynamical chiral symmetry breaking (DCSB) critical point which we study here at zero temperature and chemical potential.
@@
For a primer on the gap equation for mathematicians and fixed-point theory for physicists, see Appendix A,B. We analyze the solution set of the gap equation after the renormalization scale has been taken to infinity in the equation, see the Appendix for details. Solutions clearly also satisfy a finite renormalization scale equation for some current quark mass parameter but the reverse is not guaranteed. In the case of the gap equation, the current quark mass parameter disappears in the limit of infinite renormalization scale and so any solutions which rely on it to reach an eigenvalue of 1 also disappear with it, namely the Wigner solutions. Solutions that set their own scale as the current quark mass parameter disappears, namely the Nambu solutions, survive and are guaranteed to exist for all current quark masses $m\equiv\lim_{p\rightarrow\infty}(\log\frac{p}{\mu})^{\gamma_{m}}M(p)$ as we will prove for a popular model using the rainbow-ladder(RL) vertex[10].
@@
Existence results of non-linear integral equations $u(p)=T(u)\equiv\int_{0}^{\infty}dkK(k,p)f(u(k),k)$, often aim to establish a fixed point between two opposite behaviors, say as $\|u\|\rightarrow 0$ and $\|u\|\rightarrow\infty$ with the supremum norm via the Krasnosel’skii-Guo Cone Compression Theorem[8, 9]. The trace of the gap equation takes exactly this form after sending the renormalization scale to $\infty$[11]. Another approach called Schauder Fixed Point Theorem can be used if $T(P)\subseteq P$ for some compact input space $P$ as we will see is the case for the trace of $\not${p} times the gap equation. Coupled systems with Krasnosel’skii type conditions in both input function have been analyzed in [7] and a mixed Krasnosel’skii-Schauder theorem has been developed in [6].
@@
Taking the trace of the Euclidean space rainbow ladder gap equation, we define $u(p)\equiv r(p)\frac{M(p)}{Z(p)}\equiv r(p)B(p)$ for some continuous and positive function $r(p)$ with $\lim_{p\rightarrow\infty}r(p)\sim\log^{\delta}\frac{p}{\mu}$ [^3] We write $\lim_{p\rightarrow p_{0}}A(p)\sim B(p)$ for $\lim_{p\rightarrow p_{0}}\frac{A(p)}{B(p)}=1$. for some $0<\delta<\gamma_{m}$ because it allows bounding the gap equation by an integrable kernel and a supremum or infimum over a bounded function. [^4] For $\gamma_{m}=0$, any value of $\delta>0$ can be used. In the perturbative region, we show in the appendix that the solution is a sum of the expected $B_{+}(x)=\frac{c_{1}}{x^{\gamma_{m}}}$ with $x\equiv\log\frac{p}{\mu}$ and $B_{-}(x)=c_{2}x^{\gamma_{m}-1}e^{-2x}$. There, we prove that $c_{2}$ is always equal to a certain integral and we will show that $m\equiv c_{1}\geq 0$ is in fact free to choose for a certain guaranteed set of solutions. Consider the input space $P=\{B(x)\in C[(-\infty,\infty)]\ |\ B(x)\geq 0,\exists\gamma_{m}>0:\lim_{x\rightarrow\infty}\frac{\partial\log B}{\partial\log x}=-\gamma_{m}\}$. Since $r(p)$ is given, we can switch between $u$ and $B$ whenever convenient and $\lim_{x\rightarrow\infty}\frac{\partial\log B}{\partial\log x}=-\gamma_{m}$ implies $\lim_{x\rightarrow\infty}\frac{\partial\log T(B)}{\partial\log x}=-\gamma_{m}$ by D.1. We have
@@
with $q_{\mu}=p_{\mu}-k_{\mu}$ with $S^{3}$ the surface of the sphere that leaves $k^{2},p^{2}$ invariant. We first consider $Z(p)$ to be a known function that one can take to satisfy the gap equation if a solution exists. It should be almost everywhere continuous and converge to a perturbative tail: $Z(p)\rightarrow 1$ as $p\rightarrow\infty$. One can write $T=T^{Z}$ as the results will apply to one unchanging $Z(p)$. We extend our analysis to the coupled system of equations before the conclusion as Schauder’s theorem is required to deal with the equation for $Z(p)$. Notice $f(u,k)$ is positive and continuous and we require that $K(k,p)>0$ is continuous almost everywhere, $\int_{0}^{\infty}K(k,p)dk$ is continuous and $\mathcal{G}(q)\sim\mathcal{G}_{pert}(q)=\frac{4\pi^{2}\gamma_{m}}{q^{2}\log(\frac{q^{2}}{\mu^{2}})}$ as $q^{2}\rightarrow\infty$. [^5] QED also fulfills similar requirements if the Planck scale is taken as upper limit. We define a simplest extension $\mathcal{G}_{simp}(q)=\frac{4\pi^{2}\gamma_{m}}{\max(q^{2},\mu^{2})\log(e+\frac{q^{2}}{\mu^{2}})}$ as a toy example.
@@ 2.3 Combining the bounds
$t(p)=\frac{1}{K_{r}}\int_{0}^{\infty}dk\frac{kZ^{2}(k)}{4\pi^{4}}\int_{S^{3}}\mathcal{G}(q(k,p))t(k)\equiv\frac{1}{K_{r}}T_{c}(t)$ (8)
@@ 3.1 Krasnosel’skii-Guo Theorem for $B(p)$
In order to utilize the Krasnosel’skii-Guo Theorem, we need to augment our space with a norm. Writing $\|u\|_{\infty}\equiv\sup_{p}u(p)$, consider therefore the Banach space $P_{\gamma_{m}}=\{B(x)\in C[(-\infty,\infty)]\ |\ B(x)\geq 0,\lim_{x\rightarrow\infty}\frac{\partial\log B}{\partial\log x}=-\gamma_{m}\}$ with norm $\|u\|=\|u\|_{\infty}+\sup_{\lambda>1,p>\mu}\log^{\gamma_{m}-\delta}(\lambda)u(\lambda p)$. The purpose of the second piece will soon become apparent but it can be checked that this is indeed a Banach space. In particular, to check that the norm remains finite for $\lambda\rightarrow\infty$, we use the asymptotic bound $\lim_{k\rightarrow\infty}u(k)\sim\frac{m}{\log^{\gamma_{m}-\delta}\frac{k}{\mu}}$ so that $\lim_{\lambda\rightarrow\infty}\log^{\gamma_{m}-\delta}(\lambda)u(\lambda p)=m$. Let $P_{\Delta}$ be the $P_{\gamma_{m}}$ corresponding to $\gamma_{m}=0$ and take one choice of $B=B_{0}(p)$ with asymptotic mass $m$. Clearly, $P_{\Delta}$ is a total cone in $E$ and for any $\in P_{\gamma_{m}}, BB$$B_{0}\in P_{\Delta}$ with $\|B_{0}\|_{\infty}<\epsilon$ for some small $\epsilon$. We denote the corresponding $u_{0}(p)=B_{0}(p)r(p)$ and $T_{\Delta}(u)\equiv T(u+u_{0})-u_{0}$.
$K(\cdot,p)$ is $L^{1}$-continuous by Scheffé’s lemma which gives equicontinuity and then compactness of $T_{\Delta}$ under the $\|u\|_{\infty}$ part of the norm via the Arzelà–Ascoli theorem. For the second part of the norm, the same follows for each $\lambda$ after a variable transformation $s=p\lambda$.
@@
$=\sup_{0<n<1}\lim_{\|u\|\rightarrow\infty}\frac{n\frac{\gamma_{m}}{\delta}(1+\log^{\gamma_{m}-\delta}(\frac{\|u\|_{\infty}}{\mu})}{(1+\left(\frac{1}{1-n}\right)^{\frac{\gamma_{m}-\delta}{\delta}}\log^{\gamma_{m}-\delta}(\frac{\|u\|_{\infty}}{\mu}))}$ (10)
@@ 3.2 Schauder Theorem for $Z(p)$ and coupled equations
The gap equation for $Z(p)$ is given by Eqn.20. We require $K_{Z}(k,p)$ to be $L^{1}$-continuous. Then uniform boundedness follows by the reciprocal nature of the gap equation for $A(p)$ for an input space $\{A(p)\in C[[0,\infty)]\ |\ A(p)>A_{-}(p)>0\}$ and therefore compactness by the Arzelà–Ascoli theorem. If, as in the models in [10], the kernel $K_{Z}(k,p)$ is negative for $k<k^{*}(p)$ and positive for $k>k^{*}(p)$ for each $p$ and we define $T_{\Delta}:P^{r,R}_{\Delta,*}\rightarrow P_{\Delta}$ as before with $u(p)\equiv M(p)r(p)$ and $r(p)$ increasing, then because $M(p)$ is decreasing, the finite mass term will suppress the negative part of the kernel $K^{-}_{Z}$ more than the positive part and push $T(A)$ from $T(A_{0})$ in the direction of $A=1$ so that for $A(p)\geq\phi_{0}(p)\equiv\min[1,A_{0}(p)]$, we have $\phi_{0}(p)\leq T_{Z}(A)(p)\leq T^{+}_{Z}(\phi_{0})(p)$ for finite $M(p)$ where $T^{+}_{Z}$ is the positive part of the kernel (likewise for $T^{-}_{Z}$ so that $T_{Z}=T_{Z}^{+}+T_{Z}^{-}$). Therefore, we can use a combined Krasnosel’skii-Schauder theorem that has been derived for coupled systems[6] as long as $\sup_{r}K_{r}>1$ for $A=T_{Z}^{+}(\phi_{0})$. In fact, for this value a solution for $A(p)$ to $M=0$ can also be guaranteed by noticing that if $A_{-}(p)$ can be found such that $A(p)>A_{-}(p)$ and $A_{--}(p)\equiv((T_{Z}^{+})^{2}+T^{-}_{Z})(A_{-})>A_{-}(p)$, then $A_{--}(p)<T(A)(p)<T^{+}_{Z}(A_{-})(p)$, [^10] Sufficient condition for $A_{-}(p)$ to be a lower bound and therefore $A_{+}=T^{+}_{Z}(A_{-})$ to be an upper bound is $T^{-}_{Z}(A_{-})+T_{Z}^{+}(A_{+})>A_{-}$ for all $p$. guaranteeing a solution for $M=0$ by Schauder’s theorem and to the coupled system as long as $\sup_{r}K_{r}>1$ for $A(p)=T^{+}_{Z}(\min[1,A_{-}(p)])$. Therefore, we have $(T_{\Delta},T_{Z}):P^{r,R}_{\Delta,*}\times P_{Z}\rightarrow P_{\Delta}\times P_{Z}$ with $T_{\Delta}$ allowing us to use the Krasnosel’skii-Guo Theorem as before [^11] The version used in [6] uses a more general formulation due to Benjamin which however includes all cases where the Krasnosel’skii-Guo Theorem holds. and $P_{Z}=\{A\in C[[0,\infty)]\ |\ A_{-}(p)\leq A(p)\leq T_{Z}^{+}(A_{-})(p)\}$ a closed, convex and bounded set under sup-norm which is closed under $T_{Z}$, therefore amenable to Schauder’s Theorem. Therefore, a positive, continuous solution pair $(u(p),A(p))$ with $u(p)$ decreasing must exist for every value of $m$ by the combined theorem. We require $\sup_{r^{*}}K_{r}>1$ for $A(p)=T^{+}_{Z}(\min[1,A_{-}(p)])$ and $r^{*}(p)$ increasing so that $r(p)$ can be chosen to be increasing and $M(p)$ is decreasing.
@@ Conclusion
Considering a more general vertex, for example the Ball-Chui vertex[11, 3], we encounter serious difficulties with this approach due to the fact that a finite-difference derivative term $A^{\prime}(p)$ appears in the equation for $B(p)$ which is not constrained to have any sign or magnitude by Schauder’s Theorem. Concerning proving $M(p)\rightarrow 0^{+}$ for a more general vertex, the Krein-Rutman theorem relies on a positive kernel which is no longer guaranteed.
Part of the issue with quantum field theory is that the infinite tower of Dyson-Schwinger equations may not have a guaranteed solution space. A corollary to the results in this paper is that for Nambu solutions to the gap equation in QCD, the contractive condition at $\|u\|\rightarrow\infty$ for the Krasnosel’skii-Guo theorem is satisfied for any $m$.
@@ Appendix A Gap equation primer for mathematicians
The gap equation is a 2-component coupled integral equation for the quark mass function $M(p)$ and field renormalization function $Z(p)$ as a function of 4-momentum $p_{\mu}$ with $p\equiv\sqrt{-p^{2}}$ in terms of the gluon propagator times coupling $g^{2}(q)D_{\mu\nu}(q)$ and vertex $\Gamma_{\mu}(p,k,q)$ with $k_{\mu}$ and $q_{\mu}=p_{\mu}-k_{\mu}$ loop momenta:
[…]
with
$\Sigma(p)=iZ_{1}\frac{4}{3}\int\frac{d^{4}q}{(2\pi)^{4}}\gamma_{\mu}g^{2}(q)D^{\mu\nu}(q)S(k)\Gamma_{\nu}(k,p)$ (18)
The simplification of the rainbow-ladder vertex amounts to $\Gamma_{\mu}(p,k)=\gamma_{\mu}$ and we use Landau gauge so that $Z_{1}g^{2}(q)D_{\mu\nu}(q)=(g_{\mu\nu}-\frac{q_{\mu}{q_{\nu}}}{q^{2}})\mathcal{G}(q)$[10]. After rotating to the Euclidean metric and letting $A(p)\equiv\frac{1}{Z(p)}$ and $B(p)\equiv\frac{M(p)}{Z(p)}$, we have the operators [11, 1]